• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name K. McCormick

    Log entry time 21:35:25 on January15,2002

    Entry number 80346

    keyword=Target cooldown and vacuum problems

    The cryotarget cooldown was finished at ~9:00 pm. JP and I went into the Hall, adjusted the fan speed of loop 2 to 60 Hz, removed the theodolites and put them behind the RCS shield wall and tried to hook up some cameras (we failed since we couldn't find the right video connector to couple the camera to the BNC).

    While checking the target screens, I noticed the vacuum in the scattering chamber was 5x10-3. I called Rusty and he came in to check on the pumps. We are now in controlled access.